OpenStudy (anonymous):

There are 40 elephants. 6 are blue, 8 black, 10 red, 6 yellow and 10 green. The 40 elephants are raced. 3 of the yellow elephants and tied together whilst 4 of the blue are tied toegher, what is the probability that these 7 elephants are in top 10?

OpenStudy (anonymous):

33C3 * 7C7/ 40C10 6.44 * 10^(-6)

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

oh is it this simple.. so u dont have to form groups from the elephants that are tied together.. umm not sure.. was on test dunno ans

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

@Ganpat im not sure :x

OpenStudy (ganpat):

i can give a following try , 3 yellow elephants are tied together from group of 6 = 6C3 4 blue elephant r tied together from group of 6 = 6C4 7 from a group of 40.. So, 40C7.. (6C3 * 6C4) / 40C7 ... what say ?? i guess this might be right..

OpenStudy (anonymous):

but thats not the prob of the 7 elephants in the first 10 is it ?

OpenStudy (anonymous):

i swear we did this yesterday or was it the day before?

OpenStudy (anonymous):

different part 2

OpenStudy (anonymous):

\[\frac{\left(\begin{matrix}33 \\ 3\end{matrix}\right)}{\left(\begin{matrix}33 \\ 3\end{matrix}\right)+\left(\begin{matrix}33 \\ 7\end{matrix}\right)+\left(\begin{matrix}33 \\ 6\end{matrix}\right)+\left(\begin{matrix}33 \\ 10\end{matrix}\right) }\] is ans

OpenStudy (dumbcow):

ok how they determined the denominator is looking at the 4 possible cases -Both sets of tied elephants finish in top 10 - Only tied yellow elephants finish top 10 - Only tied blue elephants finish top 10 - Neither finish top 10 For each case, look at 33 choose n, where n is number of untied elephants who finish top 10

OpenStudy (anonymous):

yeah i think i get it .. i dont think anyone got that on the test ..

OpenStudy (dumbcow):

haha i didn't either at first, i worked backwards after seeing the answer